OpenStudy (anonymous):

Hi guys, this is a linear algebra question. Let \[V \subseteq Map(\mathbb{R},\mathbb{R}) \] be the sub-vector space of continuous functions. Let W be the sub-vector space of V spanned by the functions 1, sin x, and sin^2 x. Show that: \[1, \sin(x), \sin(x^2) \in W \] are linearly independent. I have no idea how to approach this... the only thing I know about linear independence is in relation to vectors. Not to functions.

OpenStudy (anonymous):

you have to show that if a linear combination is zero, then each coefficient must be zero

OpenStudy (anonymous):

in other words is \[a_1+a_2\sin(x)+a_3\sin^2(x)=0\] then \[a_1=a_2=a_3=0\]

OpenStudy (anonymous):

in this context, the meaning of \[a_1+a_2\sin(x)+a_3\sin^2(x)=0\]is that it is the zero function, i.e. it is zero for all values of \(x\) not just for some value of \(x\)

OpenStudy (anonymous):

at the moment, however, i totally forget how to do this, but there should be some similar example in a text. i believe you can do it by the "wronskian"

OpenStudy (loser66):

:) i have a question, sin^2x or sin (x^2) ?

OpenStudy (anonymous):

\[ f(g'h'' - h'g'') - g(f 'h'' - h'f '') +h(f 'g'' - g'f'')\]

OpenStudy (anonymous):

here is an example http://www.youtube.com/watch?v=snbmY5oQMlQ

OpenStudy (anonymous):

Sorry, took me a while to figure out how to make a pciture of it.